OpenStudy (yolo_king):

There are 4 red paint cans, 2 green paint cans, and 1 yellow paint can. What is the probability of choosing a green paint can then a yellow paint can without replacement? A) 1/21 B) 2/21 C) 7/12 D) 21/2

Mehek (mehek14):

Green \(\dfrac{2}{7}\) Yellow \(\dfrac{1}{6}\) \(\dfrac{2}{7}*\dfrac{1}{6}\)

OpenStudy (yolo_king):

2/42 = 1/21

Mehek (mehek14):

Correct